Jérémy passes a backpack, but he wants to buy a bigger one, with a capacity of 45 to 55 L. he hesitates between two models. The capacity of the model A is 2 times greater than that of the backpack it already has. As for the model B, it can hold 15 L more than its walking backpack. If we added 18L to the capacity of the model A and 28 L to the capacity of model B, both models would have exactly the same capacity, which of the two models should jérémy choose?
Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

A=2x
B=x+15
so 2x+18=x+43
x=25 liters
so model A would be 50 liters and in his range, whereas model B would be 40 liters and not within his range.

Answer by greenestamps(13200)



I can't imagine that you could POSSIBLY have looked at your post before you posted it....

And I can't imagine, with the way the problem is stated, that the other tutor purported to understand the given information....

Perhaps with the person's name being Jérémy, the original problem was in Hungarian and was translated poorly.

Or perhaps the problem was dictated orally and the voice recognition program did a poor job of translating it into words....

--------------------------------------------------------------

"Jérémy passes a backpack, but he wants to buy a bigger one..."

From the context, he has a backpack but want to get a bigger one; maybe "passes" is supposed to be "possesses"?

--------------------------------------------------------------

"The capacity of the model A is 2 times greater than that of the backpack it already has."

The "it" to replace "Jérémy" suggests that maybe the problem was indeed translated poorly from its original language.

And even disregarding the other sloppy language, there is a fatal flaw in the statement of the problem in this sentence. Saying that the capacity of model A is "2 times greater than..." is probably intended to mean 2 times AS GREAT AS; but 2 times GREATER THAN x means x, plus 2 more times x, which is 3x, or 3 times AS GREAT. So we don't know whether that statement is really supposed to mean the capacity if 2 times as great or 3 times as great.

--------------------------------------------------------------

"As for the model B, it can hold 15 L more than its walking backpack."

"its walking backpack"???!

There is that "it" again that presumably has replaced "Jérémy"; but what on earth is a walking backpack?

--------------------------------------------------------------

For goodness sake, take the time to make your posts in such a way that we know what the question is. You are wasting our time (and yours) by posting garbage like this.